1. General status updates/Schedule/Purchases review.
2. Status of DTC to CRV ROC interface.
3. Status of EPICS progress.
4. Status of core ROC firmware testing.
Eric -- we deployed new otsdaq/artdaq release. Been working a lot with Giani to get simulation/trigger filters working. Giani converging, currently cant stop state machine.
John -- Fixed last timing card so now we have 88 of 88 functional (+5 v2 for STM functional). Wants to disassemble patch panes. Should take an hour or so (Vivian is good with it). DTCs are on site.
Vivian -- Worked with Monica last week, tried to incorporate ROC core into Tracker ROC and using Monica’s DDR interface. Things do not seem robust. Would like to proceed with artdaq chain in same vein as Giani.
Stefano -- Simona is ready to try dqm again. Giani is currently using the latest offline, so it would be a good time.
Antonio -- Last week improved server side PV subscribing/unsubscribing; web client reads value from server cache; Now adding features to web client window; adding menus, etc. Gave demonstration monday of EPICS interface control dashboard (to Glenn and Michael Hedges and Stefan).
Stefan -- Trying to repeat when Antonio is doing and test features. Worked with Oliver on STM FPGA things - connected the ADC card. Power cycling changed the firmware. Communicating with Erdem - will check in code. Oliver is going to try algorithm on HLS, defined interface with Erdem. Oliver returning Saturday.
Oliver -- has link to Erdem’s Redmine repo. Likely want Vivado/HLS installed on daq cluster
Rick -- debugging hardware event building ..fixing lots of bugs. Sten and Lei came over.. They need to update the CRV box we have. We are going to try SMA interface for clock/marker.
Karen -- heard about SMA at CRV meeting, thought there was a CRV processing issue
Bertrand -- Worked on improving a module in track trigger. 3.4ms on daq07 20 cores→ one core is 2.8 … think we saved 0.8 ms with new changes. We may be approaching a plateau. Need to start considering unpacking.
Greg -- Are racks lockable? Yes. Keys are currently in PREP. Took a look at Error/accounting doc again - would like a clear description of overall accounting plan (Bertrand suggested companion document ⇒ what specifically are we saving and what is sourcing that information). The existing document descibes technically what should be possible.
Simona -- interested in Giani’s progress.
Ryan -- DTC’s are on site now, not in hand, but accounting accepted them already. ROC core functionality is in disarray with no one working on it. Demo card needs to be stabilized.